# Hallucination Detection in LLMs with Topological Divergence on Attention Graphs

### v2

@@ 6 Conclusion
This paper introduces TOHA, a novel hallucination detection method based on the topological divergence of attention maps. At its core, TOHA leverages our key observation that specific attention heads exhibit consistent patterns during hallucinations — regardless of the dataset. TOHA computes hallucination scores by averaging the topological divergences from these heads, and we formally prove several stability properties to ensure these scores are reliable. Additionally, we explored the behaviour of “hallucination-aware” heads, discovering that the attention to <s> token plays an important role in their discriminative ability. This importance of <s> aligns well with prior work Barbero et al. (2025).
Extensive experiments show that TOHA is a robust alternative to existing approaches, matching or surpassing state-of-the-art baselines like SelfCheckGPT Manakul et al. (2024). Notably, TOHA is significantly more efficient, operating up to an order of magnitude faster than methods of comparable quality. We further validate TOHA’s transferability, demonstrating its robustness to shifts in data distribution — a critical advantage for real-world deployment, where LLM inputs are far more diverse and complex than benchmark examples.
In summary, TOHA delivers state-of-the-art detection performance while combining efficiency and solid generalizability, making it particularly suited for practical applications.
@@
RAG scenario. While TOHA operates effectively in the RAG scenario under the assumption that the provided context contains the correct answer, we recognize that this condition may not always hold in real-world applications. This limitation points to an important direction for future research, where the method could be extended to handle cases of incomplete or unreliable context knowledge.
@@
Appendix A Topological data analysis: background
@@
$C_{\bullet}:\cdots\rightarrow C_{k+1}\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}\cdots,\,$
@@
[…]
@@
Appendix B MTop-Div on graphs properties
Basic properties of MTop-Div for attention graphs.
Now we consider specific properties for our adaptation of $\operatorname{MTop-Div}(R,P)$.
Proposition B.1.
The following holds for any attention graph $G$ and its complementary vertex subsets $P,R\subset~{}V_{G}$.
  • $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
  • Let the natural norm on the cross-barcodes be defined as follows:
    $\lVert\operatorname{Cross-Barcode}_{0}\rVert_{B}=\\ \max_{[b_{j},d_{j}]\in\operatorname{Cross-Barcode}_{0}}(d_{j}-b_{j}).$ (3)
    The norm of $\operatorname{Cross-Barcode}_{0}(R,P)$ lays in the interval [0, 1]01[0,\,1][ 0 , 1 ]:
    $0\leq\lVert\operatorname{Cross-Barcode}_{0}(R,P)\rVert_{B}\leq 1$ (4)
@@ Appendix B MTop-Div on graphs properties
    $0\leq\operatorname{MTop-Div}(R,P)\leq|R|.$ (5)
The second and third statements are immediately obtained from the properties of an attention matrix: all its weights lie between 00 and $1$.
The following property formalizes the intuition behind our metric — it measures the strength of the response’s connection to the prompt through multiscale topological features of the attention graph.
Proposition B.2.
@@
Proof of Proposition B.2.
We have to check the definition of the exact sequence: $\text{Ker}(r_{i})=\text{Im}(r_{i+1})$. For a pair $r_{0},\,r_{1}$, it is equivalent to the surjectvity of $r_{1}$. The $H_{0}$ homology group of a graph corresponds to the connected components of the graph. The set of edges $E_{(G,w)}^{\leq\alpha}=\{e\in E_{G}|w_{e}\leq\alpha\}$ is always a subset in the analogous set of the weighted graph $(G,w_{(R\cup P)/P})$ with all weight edges between $P$ vertices set to zero. Therefore, the map $r_{1}$ between their connected components is surjective. Similarly, the kernel of the map $r_{1}$ is spanned by the differences of two connected components, which are merged after adding some of the edges between $P$ vertices, and any such difference lies in the image of the map $r_{2}$. Also, any two vertices from $P$ belong to the same connected component in the graph $(G,w_{(R\cup P)/P}\leq\alpha)$, hence the image of $r_{2}$ is in the kernel of $r_{1}$. Therefore, the considered sequence is exact indeed. ∎
Proof of Proposition 3.1.
1. The $0-$th Cross-Barcode coincides with the set of edges in the minimal spanning tree of the weighted graph $G$ with all the weights within $P$ vertex subset equal zero. Excluding the zero weight edges, this edge set coincides with the minimal spanning forest attaching the vertex set $R$ to $P$ vertices. ∎
2. Denote by $\operatorname{MSF}(R,P)$ the minimum spanning forest attaching $R$ to $P$. Note that we have properties B.1, so
@@
There are two possibilities: 1) after a change, all MSF edges remain the same, or 2) some edges are replaced with other edges. In the first case, it is obvious that the total sum of edge weights changes by no more than $\delta=\varepsilon\cdot\#\text{edges}(\operatorname{MSF}(R,P))=\varepsilon\cdot|R|$. Consider the second case. Denote by $\operatorname{MSF}_{\mathrm{prev}}$ the original MSF, by $\operatorname{MSF}_{\mathrm{new}}$ — the MSF after the change; let $w$ be the edge weight function before the change, $\hat{w}$ — after the change. The following inequalities hold:
@@
3. Follows obviously from the MSF formula for $\operatorname{MTop-Div}(R,P)$ and attention map properties.
Appendix C Datasets
@@
Prompt number 1 2 3 4 5
CoQA Accuracy ($\uparrow$) 0.809 ± 0.017 0.861 ± 0.015 0.742 ± 0.003 0.795 ± 0.009 0.831 ± 0.025
Precision ($\uparrow$) 0.849 ± 0.021 0.911 ± 0.007 0.771 ± 0.003 0.828 ± 0.011 0.860 ± 0.012
Recall ($\uparrow$) 0.871 ± 0.004 0.877 ± 0.019 0.877 ± 0.013 0.877 ± 0.005 0.893 ± 0.027
SQuAD Accuracy ($\uparrow$) 0.831 ± 0.003 0.857 ± 0.018 0.857 ± 0.008 0.872 ± 0.003 0.854 ± 0.007
Precision ($\uparrow$) 0.813 ± 0.002 0.831 ± 0.028 0.845 ± 0.021 0.850 ± 0.011 0.847 ± 0.007
Recall ($\uparrow$) 0.796 ± 0.008 0.839 ± 0.010 0.823 ± 0.023 0.858 ± 0.018 0.813 ± 0.017
Average Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$)
CoQA 0.808 0.844 0.879
SQuAD 0.854 0.837 0.826
Table 7: Classification metrics of GPT-4o annotation for CoQA and SQuAD with human labels considered actual annotation. The top table shows metric scores for different variants of prompts used. The bottom table shows the metric scores averaged across all prompt variants.
C.1 Data Generation & Annotation

### v3

@@ 5 Conclusion
This paper introduces TOHA, a novel hallucination detection method based on the topological structure of attention maps. Central to TOHA is the $\operatorname{MTop-Div}_{G}(R,P)$, our adaptation of Manifold Topology Divergence Barannikov et al. (2021) for the graph setting, which, as we demonstrate, serves as a measure of the response novelty in relation to the prompt. This property makes the proposed divergence well-suited for hallucination detection in RAG scenarios.
By analyzing the divergence values, we identified a subset of “hallucination-aware” attention heads that reliably distinguish hallucinated and grounded samples, irrespective of the dataset. TOHA computes the final hallucination scores by averaging the topological divergences from these heads. Further investigation reveals that some of these heads are associated with copying behavior, which aligns with prior work Sun et al. (2025).
Extensive experiments show that TOHA is a robust alternative to existing approaches, matching or surpassing state-of-the-art baselines. Moreover, our method is both data- and compute-efficient: just $50$ annotated samples suffice for reliable detection, and inference runs several times faster than comparable methods of similar quality. Crucially, we validate TOHA’s transferability, demonstrating its robustness to shifts in data distribution — a key advantage for real-world deployment, where LLM inputs are far more diverse and complex than benchmark examples.
In summary, TOHA delivers state-of-the-art detection performance while combining efficiency and solid generalizability, making it particularly well-suited for practical applications.
@@
Appendix A Related Works
Hallucination detection methods.
The problem of hallucinations in LLMs has attracted significant attention recently Zhang et al. (2023); Huang et al. (2023); Wang et al. (2024). Consistency-based methods  Manakul et al. (2024); Chen et al. (2024); Qiu and Miikkulainen (2024); Nikitin et al. (2024); Hou et al. (2025) that use the diversity of multiple LLM responses as a hallucination score offer robust detection but impose significant computational overhead. Surface-level techniques like perplexity and logit entropy Fadeeva et al. (2024); Malinin and Gales (2021) analyze model confidence directly from output distributions — efficient but limited in detection capability as they neglect the model’s rich internal representations Gekhman et al. (2025). Hidden states-based classifiers Azaria and Mitchell (2023); Sky et al. (2024); Zhou et al. (2025) require extensive annotated datasets, which are scarce in the public domain Zhang et al. (2023), and suffer from poor task transferability Sky et al. (2024). While HaloScope Du et al. (2024) partially addresses the data scarcity issue by leveraging unlabeled “in-the-wild” generations, it still demands a large volume of unannotated model outputs to achieve strong performance.
Attention map-based methods represent a promising yet underdeveloped direction. Current techniques either (i) rely on large annotated datasets (e.g., Lookback Lens Chuang et al. (2024), LapEigVals Binkowski et al. (2025)), (ii) exploit only simplistic attention graph properties like self-loop weights (LLM-Check Sriramanan et al. (2024)), or (iii) disregard attention map geometry entirely, using mechanistic scores instead (ReDEEP Sun et al. (2025)). This leaves a critical research gap: training-free methods that fully leverage the rich structural information encoded in attention graphs remain underexplored.
Topological Data Analysis (TDA) in NLP.
Topological Data Analysis (TDA) is a mathematical framework for extracting multi-scale structural patterns from data using principles from topology and computational geometry Chazal and Michel (2017); Hensel et al. (2021). Recent years have seen growing interest in applying TDA to natural language processing (NLP) tasks to study textual structural properties Uchendu and Le (2024). For example,  Tulchinskii et al. (2024) leveraged persistent homology to estimate the intrinsic dimensionality of CLS embeddings for detecting machine-generated text. Other work has demonstrated the utility of topological features derived from transformer attention matrices — treated as weighted graphs — for diverse NLP applications. These include uncertainty quantification Kostenok et al. (2023) and grammatical acceptability classification Proskurina et al. (2023), where topological features extracted from the attention graphs were used as input to train auxiliary classifiers.
Appendix B Additional Experiment Results
B.1 Multiple generations-based methods
To provide a complete comparison, we considered an implementation of consistency-based methods with $20$ additional generations. The results are demonstrated in Tables 5-6. We can see that TOHA still remains superior to the baselines, achieving top performance in most experiments.
Table 5: ROC AUC $(\uparrow)$ of multiple generations-based methods (with 20 additional samples) and TOHA. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single MS MARCO CNN/DM + CoQA SQuAD XSum
generation Recent News
Mistral-7B
SelfCheckGPT [1] 0.67 $\pm$ 0.03 0.59 $\pm$ 0.04 0.93 $\pm$ 0.01 0.83 $\pm$ 0.02 0.71 $\pm$ 0.04
Semantic entropy [2] 0.53 $\pm$ 0.03 0.52 $\pm$ 0.03 0.86 $\pm$ 0.01 0.74 $\pm$ 0.01 0.63 $\pm$ 0.02
EigenScore [3] 0.49 $\pm$ 0.03 0.53 $\pm$ 0.05 0.78 $\pm$ 0.02 0.77 $\pm$ 0.04 0.59 $\pm$ 0.05
TOHA (ours) 0.76 $\pm$ 0.04 0.60 $\pm$ 0.09 0.89 $\pm$ 0.01 0.96 $\pm$ 0.01 0.66 $\pm$ 0.05
LLama-2-7B
SelfCheckGPT [1] 0.60 $\pm$ 0.04 0.60 $\pm$ 0.03 0.77 $\pm$ 0.02 0.78 $\pm$ 0.02 0.67 $\pm$ 0.04
Semantic entropy [2] 0.56 $\pm$ 0.03 0.49 $\pm$ 0.03 0.79 $\pm$ 0.01 0.77 $\pm$ 0.02 0.63 $\pm$ 0.04
EigenScore [3] 0.57 $\pm$ 0.04 0.52 $\pm$ 0.06 0.61 $\pm$ 0.03 0.78 $\pm$ 0.02 0.65 $\pm$ 0.03
TOHA (ours) 0.65 $\pm$ 0.02 0.56 $\pm$ 0.02 0.90 $\pm$ 0.01 0.87 $\pm$ 0.04 0.68 $\pm$ 0.05
LLaMA-2-13B
SelfCheckGPT [1] 0.61 $\pm$ 0.05 0.60 $\pm$ 0.06 0.88 $\pm$ 0.02 0.75 $\pm$ 0.04 0.61 $\pm$ 0.04
Semantic entropy [2] 0.60 $\pm$ 0.03 0.52 $\pm$ 0.03 0.77 $\pm$ 0.04 0.70 $\pm$ 0.02 0.62 $\pm$ 0.03
EigenScore [3] 0.58 $\pm$ 0.04 0.48 $\pm$ 0.05 0.59 $\pm$ 0.03 0.60 $\pm$ 0.02 0.54 $\pm$ 0.05
TOHA (ours) 0.67 $\pm$ 0.04 0.56 $\pm$ 0.05 0.92 $\pm$ 0.02 0.88 $\pm$ 0.05 0.66 $\pm$ 0.03
Table 6: ROC AUC $(\uparrow)$ of multiple generations-based methods (with 20 additional samples) and TOHA. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single gen. CoQA SQuAD XSum
LLaMA-3.1-8B
SelfCheckGPT [1] 0.95 $\pm$ 0.01 0.78 $\pm$ 0.03 0.75 $\pm$ 0.03
Semantic entropy [2] 0.82 $\pm$ 0.03 0.54 $\pm$ 0.06 0.46 $\pm$ 0.05
EigenScore [3] 0.84 $\pm$ 0.01 0.56 $\pm$ 0.03 0.48 $\pm$ 0.03
TOHA (ours) 0.84 $\pm$ 0.01 0.87 $\pm$ 0.03 0.65 $\pm$ 0.05
Qwen2.5-7B
SelfCheckGPT [1] 0.75 $\pm$ 0.02 0.77 $\pm$ 0.02 0.72 $\pm$ 0.03
Semantic entropy [2] 0.83 $\pm$ 0.05 0.58 $\pm$ 0.04 0.69 $\pm$ 0.05
EigenScore [3] 0.83 $\pm$ 0.05 0.56 $\pm$ 0.03 0.48 $\pm$ 0.03
TOHA (ours) 0.79 $\pm$ 0.05 0.77 $\pm$ 0.02 0.69 $\pm$ 0.03
B.2 Comparison to ReDEEP
For a comprehensive comparison with ReDEEP Sun et al. (2025), which was evaluated on the entire RAGTruth dataset in the original paper, we conducted this evaluation of TOHA under the same conditions. As shown in Table 7, TOHA not only outperforms ReDEEP but also demonstrates greater robustness to data distribution shifts, exhibiting less significant performance degradation on separate benchmarks of RAGTruth (Tables 1-2).
Method Llama-2-7B Llama-2-13B
ReDEEP [8] 0.68 $\pm$ 0.02 0.77 $\pm$ 0.01
TOHA (ours) 0.70 $\pm$ 0.04 0.80 $\pm$ 0.02
Table 7: Performance comparison between ReDEEP and TOHA on the entire RAGTruth dataset.
B.3 Alternative attention map-based features for hallucination detection
Features MS MARCO CoQA
Mistral-7B
Standard topological 0.67 0.69
Sparsity ratio 0.66 0.7
Entropy 0.75 0.77
Wasserstein 0.77 0.73
Spectral norm 0.73 0.72
Attention to <s> 0.65 0.61
MTop-Div 0.86 0.98
LLaMA-2-7B
Standard topological 0.69 0.7
Sparsity ratio 0.49 0.61
Entropy 0.38 0.68
Wasserstein 0.73 0.6
Spectral norm 0.49 0.64
Attention to <s> 0.62 0.64
MTop-Div 0.75 0.96
During preliminary experiments for an attention map-based hallucination detector, we evaluated a range of topological and traditional features. As standard topological features, we employed barcode-based features, such as the sum of bar lengths in persistence diagrams, and naive topological features, including the average vertex degree in attention graphs Kushnareva et al. (2021). As for traditional attention-based features, we used sparsity ratio, attention entropy, and spectral norm Kobayashi et al. (2020); Vig and Belinkov (2019); Ji et al. (2021). We also considered Wasserstein distances between the persistent diagrams Edelsbrunner and Harer (2010) of the prompt and response subgraphs as an alternative way to measure their similarity. Finally, we analyzed the average attention to the first token as we discovered that hallucination-aware heads often attend to in presence of hallucinations (see Section 4).
To identify the most informative features for hallucination detection, we trained $L1$-regularized supervised classifiers on concatenated features from all layers and heads and compared them to classifiers trained only on $\operatorname{MTop-Div}_{G}(R,P)$ values under the same conditions. The results are presented in Table 8.
While the classifier based on $\operatorname{MTop-Div}_{G}(R,P)$ values significantly outperforms alternative approaches, computing these values across all layers and attention heads is highly computationally expensive. To address this, we developed TOHA — a more efficient alternative that aggregates divergence values from only a subset of the “hallucination-aware” attention heads.
B.4 Other metrics for the head selection procedure
Additionally, we investigated alternative attention-map-based scores — including entropy, spectral norm, and the Wasserstein distance between the persistent diagrams of prompts and responses — for selecting specialized attention heads. Following the pipeline of the Algorithm 1, we computed the average distances between hallucinated and grounded samples using alternative scores. The results, presented in Figure 10, reveal that the heads that separate samples best for MS MARCO do not generalize to the CoQA dataset. This suggests that our proposed $\operatorname{MTop-Div}_{G}(R,P)$ metric is better suited for the task compared to existing solutions.
B.5 Performance analysis
Additionally, we analysed TOHA’s performance for extremely short responses consisting of one word. The results are presented in Figure 7. While its performance drops — as expected when the response graph degenerates to a single vertex — it remains non-random. This indicates that TOHA retains some predictive power even for these challenging, out-of-design cases.
As for the TOHA’s performance on longer responses, the results for the MS MARCO, CNN/DM and XSum datasets demonstrate that as the response length grows, the hallucination signals become less distinguishable for all the considered methods, as hallucinated tokens comprise a very small part of the response. However, TOHA remains superior to baselines, confirming that our topology-based approach is more effective than existing methods for complex long-form responses.
Figure 7: Comparison of ROC-AUC scores ($\uparrow$) for single-word versus multi-word model responses. Dataset: SQuAD.
B.6 P-values for the transferability experiment
As we mentioned, for CNN/DM and Xsum dataset, the TOHA’s performance in transfer settings fall within the standard deviation of the method. To confirm this, we calculated the p-vales of the corresponding T-test for means of two independent samples. The results, provided in Table 9, support our claim: all p-values are significantly greater than $0.05$.
Probe set CNN/DM XSum
MS MARCO 0.272 1.0
CNN/DM 1.0 0.746
CoQA 0.499 0.134
SQuAD 0.373 0.346
XSum 0.515 1.0
Table 9: P-values of the T-test for the transferability experiment.
Appendix C Topological Data Analysis: Background
@@
$C_{\bullet}:\cdots\rightarrow C_{k+1}\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}\cdots,\,\$
@@
As the threshold $\varepsilon$ increases, new topological features (e.g., connected components, holes) can appear and disappear. The persistent homology tool tracks the dynamics of these topological features. Formally, the $k$-th persistent homology of $S$ is the pair of sets of vector spaces $\{H_{k}(S_{i})\,|\,0\leq i\leq n\}$ and maps $f_{ij}$, where $f_{ij}:H_{k}(S_{i})\rightarrow H_{k}(S_{j})$ is a map induced by the embedding $S_{i}\subseteq S_{j}$. Each persistent homology class in this sequence is “born” at some $S_{i}$ and “dies” at some $S_{j}$ or never dies Barannikov (1994). This birth-death process of a basic set of independent topological features can be visualized as the set of intervals $\varepsilon_{\mathrm{birth}},\varepsilon_{\mathrm{death}}]$ called barcode (see Figure 8). The features with 0 lifespans are typically excluded. The horizontal axis is a sequence of thresholds $\varepsilon$, and each horizontal bar corresponds to a single feature. We begin with $|X|=m$ connected components (all of them are “born”), and as $\varepsilon$ increases, their pairs are merged (each merge corresponds to a “death” of a feature). The $0$-th barcode construction procedure is equivalent to Kruskal’s algorithm for minimum spanning tree (MST), the bars in the barcode correspond to the edges in the MST of $X$ Tulchinskii et al. (2023).
@@
Appendix D Properties of $\operatorname{MTop-Div}_{G}(R,P)$
Proof of Proposition 3.1.
The $0$-th Cross-Barcode coincides with the set of edges in the minimal spanning tree of the weighted graph $G$ with all the weights within $P$ vertex subset equal zero. Excluding the zero weight edges, this edge set coincides with the minimal spanning forest attaching the vertex set $R$ to $P$ vertices. ∎
Other properties of MTop-Div for attention graphs.
Here we present some other properties of our proposed $\operatorname{MTop-Div}_{G}(R,P)$.
Proposition D.1.
@@ Appendix D Properties of $\operatorname{MTop-Div}_{G}(R,P)$
    $0\leq\operatorname{MTop-Div}_{G}(R,P)\leq|R|.$ (4)
  • (Stability.) If the weights of $G$ change by no more than $\varepsilon$, then the corresponding $\operatorname{MTop-Div}(R,P)$ changes by no more than $\delta=\varepsilon|R|$.
@@
  • (Connection with hallucinations.) The normalized divergence value $\frac{1}{|R|}\operatorname{MTop-Div}(R,P)=0$ iff the MSF attaches every response token to a prompt token by a subtree with attention weights =1=1.
Proof of Proposition D.1.
1. This property is immediately obtained from the properties of an attention map: all its weights lie between 0 and $1$. ∎
2. Denote by $\operatorname{MSF}(R,P)$ the minimum spanning forest attaching $R$ to $P$. Note that we have properties D.1, so
@@
There are two possibilities: 1) after a change, all MSF edges remain the same, or 2) some edges are replaced with other edges. In the first case, it is obvious that the total sum of edge weights changes by no more than
$\delta=\varepsilon\cdot\#\text{edges}(\operatorname{MSF}(R,P))=\varepsilon\cdot|R|.$
Consider the second case. Denote by $\operatorname{MSF}_{\mathrm{prev}}$ the original MSF, by $\operatorname{MSF}_{\mathrm{new}}$ — the MSF after the change; let $w$ be the edge weight function before the change, $\hat{w}$ — after the change. The following inequalities hold:
@@
3. We have to check the definition of the exact sequence: $\text{Ker}(r_{i})=\text{Im}(r_{i+1})$. For a pair $r_{0},\,r_{1}$, it is equivalent to the surjectvity of $r_{1}$. The $H_{0}$ homology group of a graph corresponds to the connected components of the graph. The set of edges $E_{(G,w)}^{\leq\alpha}=\{e\in E_{G}|w_{e}\leq\alpha\}$ is always a subset in the analogous set of the weighted graph $(G,w_{(R\cup P)/P})$ with all weight edges between $P$ vertices set to zero. Therefore, the map $r_{1}$ between their connected components is surjective. Similarly, the kernel of the map $r_{1}$ is spanned by the differences of two connected components, which are merged after adding some of the edges between $P$ vertices, and any such difference lies in the image of the map $r_{2}$. Also, any two vertices from $P$ belong to the same connected component in the graph $(G,w_{(R\cup P)/P}\leq\alpha)$, hence the image of $r_{2}$ is in the kernel of $r_{1}$. Therefore, the considered sequence is exact indeed. ∎4. Follows obviously from the MSF formula for $\operatorname{MTop-Div}(R,P)$ and attention map properties. ∎
Intuition behind MTop-Div properties and hallucination detection.
The stability property guarantees that similar attention patterns yield similar hallucination scores, making the metric’s behavior consistent and predictable. The exact sequence property formalizes geometric intuition behind our metric — it measures the strength of the response’s connection to the prompt through multiscale topological features of the attention graph. In the last property, we present an “ideal” case: if a model “knows what to look at” — each token in the response attends to some token in the prompt with an attention weight equal to 1 — $\operatorname{MTop-Div}$ would be equal to 0, indicating zero uncertainty.
Appendix E Datasets
@@
Prompt number 1 2 3 4 5 Average
CoQA Accuracy ($\uparrow$) 0.809 ± 0.017 0.861 ± 0.015 0.742 ± 0.003 0.795 ± 0.009 0.831 ± 0.025 0.808
Precision ($\uparrow$) 0.849 ± 0.021 0.911 ± 0.007 0.771 ± 0.003 0.828 ± 0.011 0.860 ± 0.012 0.844
Recall ($\uparrow$) 0.871 ± 0.004 0.877 ± 0.019 0.877 ± 0.013 0.877 ± 0.005 0.893 ± 0.027 0.879
SQuAD Accuracy ($\uparrow$) 0.831 ± 0.003 0.857 ± 0.018 0.857 ± 0.008 0.872 ± 0.003 0.854 ± 0.007 0.854
Precision ($\uparrow$) 0.813 ± 0.002 0.831 ± 0.028 0.845 ± 0.021 0.850 ± 0.011 0.847 ± 0.007 0.837
Recall ($\uparrow$) 0.796 ± 0.008 0.839 ± 0.010 0.823 ± 0.023 0.858 ± 0.018 0.813 ± 0.017 0.826
Table 13: Classification metrics of GPT-4o annotation for CoQA and SQuAD with human labels considered as true. The table shows metric scores for different variants of prompts used and the score value averaged across all prompt variants.
E.1 Data Generation & Annotation